OpenStudy (anonymous):

You can throw an unlimited number of darts at the dartboard. Some total scores are impossible to obtain, such as all numbers less than 5 as well as 6, 8, and 9. What is the highest whole number score that is impossible to obtain?

OpenStudy (anonymous):

Here's the choices 31 18 23 29

OpenStudy (anonymous):

I think 23

OpenStudy (calculusxy):

Yes you're correct according to the Frobenius number of (a,b)

OpenStudy (welshfella):

I dont get this question

OpenStudy (calculusxy):

This is basically with the formula from Frobenius number of (a,b) \[(a,b) \] is \[(a \times b) - a - b \]

OpenStudy (calculusxy):

Thus, if we have a = 5 and b = 7 we will get \[(5 \times 7) - 5 - 7 = 35 - 5 - 7 =20 - 7 = 23 \]